• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 11:12:33 on August 5,2005

    Entry number 150792

    Followups:

    keyword=Q2 measurements

    Spectr. DAQ Run 2362, 2363, 2364 check DAQ, and spot++
    09:54 HAP. run 4134 finished, we switch to spectr. DAQ.
    HAPPEX det. goes to 1900(L), 1800(R).

    Beam current 0.6 uA

    Rates:
    T1: 495 kHz
    T2: 67.3 kHz
    T3: 470 kHz
    T4: 88.7 kHz
    T5 = T6 = 0
    T7 = 518 kHz
    T8 = 557 kHz

    Run 2365. Scint on, VDCs still off.
    Turning VDCs on.
    Run 2366. Looks good. Run ended at 620K events because beam
    was lost.
    Run 2367, repeat after the beam came back.

    I have adjusted the "db_run.dat" to have a first guess at the
    input parameters. Ebeam = 2.750 - dE/dx(3 MeV) = 2.747 GeV.
    Angles taken to be 6 degr. for now. The resulting Qsq plots
    are shown in fig 1, for each trigger. A cut on HAPPEX detector
    ADC requires that some energy be in the detector; also cut on
    number of tracks = 1 (the latter isn't too important).
    No other cleanup cuts, for now.
    The Q2 is nearly the same for each of the 3 triggers.




    FIGURE 1